V=(x2-1)/(k2+xk)(1/(1-(1/k))-1)((x-xk3-k4+k)/(1-x2))


1
Expert's answer
2020-08-04T17:07:38-0400

V = {(x2-1)/(k2+xk)} {1/(1-(1/k)-1)} {(x-xk3-k4+k)/(1-x2)}

∴ {(x2-1)/(1-x2) = -1

V = {(-1)/(k(x+k))} {1/(k-1)} {(x+k) - xk3 - k4 }

V = {(-1)/(k(x+k))} {1/(k-1)} {1(x+k) - k3(x+k)}

V = {(-1)/(k(x+k))} {1/(k-1)} {(x+k)(1-k3)}

"\therefore" many terms are cancel out after multiplication

V = -(1-k3)/{k(k-1)}

V = (k3-1)/(k2-k)=(k2+k+1)/k